Day: June 10, 2020

MFF Equipment Status

Instrument Status Status Notes
6″ Sputter deposition  Ready Air leak on the gate valve pneumatic cylinder

Cr – 100 W 10 nm/min Load: 3.49 Tune: 1.66
Au – 100 W 30 nm/min Load: 4.0 Tune: 1.46
Al – 100 W 20 nm/min Load: 4.83 Tune: 0.44
SiO2 – 100 W 5 nm/min Load:  Tune:

8″ Sputter deposition Ready Issues – Need to fix the hoist or take weight off the system
Denton e-beam deposition  Ready
Fredrick e-beam deposition  Unavailable Cryo pump and compressor needs a rebuild
Parylene Coater  Ready
TRION  Ready Rough pump needs to be rebuild, but it is operational with the CAIBE rough pump which gives a better vacuum
CAIBE  Unavailable Need to test with water cold stage and determine the beam current
March  Ready Cleaning: 200 W, 100 sccm O2, 300 secs
PDMS bonding:
40-50 W (maybe hard to get reflective power below 10 W), 100 sccm O2, 60 secs, after plasma treatment place the PDMS and glass together and heat treat at 80oC for 60 mins
Stripping carbonized SU-8: 300 W, 100 sccm O2, 300 secs (may take longer, depending on thickness)
Ellipsometer  Ready
FTIR  Ready Issues – Calibration
Curvature System  Ready Issues – Calibration
Oxidation Furnace  Ready
Diffusion Furnace  Ready
2″ Furnace  Ready
RTP  Unavailable TC wafer needs replacing and spot welded onto the connectors
EasyNano CVD CNT Ready Issue: Need to replace the pressure switch in the gas cabinet as it is flaky
Dicing Saw  Ready
EV620 Mask Aligner  Ready Turning off the lamp to preserve the life timeIssues – Logging sometimes fails as the optics stage gets stuck, the y motor also losses firmware if system is off too long
Wet bench – Chemical processing  Ready
Vacuum oven  Ready
Hotplates  Ready
Spinner  Ready
Wire bonder  Ready
Filmertics 3D Profiler Ready
Cleanroom Optical Microscopy  Ready Issues – Calibrate the CCD capture measurements
Probe station  Ready Issues – Need to install new probes from eBay
4200A – SCS Parameter Analyzer  Ready
Four point conductivity  Unavailable
Atomic Layer Deposition Ready 20190401 – New TDMAH installed and Hf deposition tuned to give a 4″ uniform deposition, growth rate is too fast per cycle, need to calibrate for the ALD region

20181205 – Milled a trench in the sleeve, put in the old ceramic pins, put in metal pins to stop the sleeve from moving

20181015 – Redo the N2 to be SS instead of plastic

201805023 – set up the process to represent an ALD valve that has constant Ar carrier flow

20180601 – Installed Ferrocene iron ALD source

20180814 – changed the precursor setup – Ferrocene is not producing a vapor pressure?

20190102 – Replaced the manual shut off valve on the ferrocene with an all metal valve – finally do not see F – not sure if it was the valve. Disconnected all the lines to the shower head that is not used.

20190219 – Installed the TDMAH to test uniformity and optimize the system

Big blue wet benches Useable 20181015 – Connected the 3-phase supply to the left side
Connected the aspirator, ran a new waste line to the current acid neutralization system. Need to install an automatic water shut off valve to the house water as the shut off requires a 10′ ladder to get to.